OpenStudy (anonymous):

How do you solve the derivative of (2x^2+4x-3)/sqrt(2x-1) ? Could you please write out the steps to get to the result?

OpenStudy (anonymous):

(6 x^2-1)/(2 x-1)^(3/2)

OpenStudy (anonymous):

Possible derivation: d/dx((2 x^2+4 x-3)/sqrt(2 x-1)) Use the product rule, d/dx(u v) = v ( du)/( dx)+u ( dv)/( dx), where u = 1/sqrt(2 x-1) and v = 2 x^2+4 x-3: (d/dx(2 x^2+4 x-3))/sqrt(2 x-1)+(2 x^2+4 x-3) (d/dx(1/sqrt(2 x-1))) Using the chain rule, d/dx(1/sqrt(2 x-1)) = -(( du)/( dx))/(2 u^(3/2)), where u = 2 x-1 and ( d1/sqrt(u))/( du) = -1/(2 u^(3/2)): (2 x^2+4 x-3) (-(d/dx(2 x-1))/(2 (2 x-1)^(3/2)))+(d/dx(2 x^2+4 x-3))/sqrt(2 x-1) Differentiate the sum term by term and factor out constants: (2 (d/dx(x^2))+d/dx(-3)+4 (d/dx(x)))/sqrt(2 x-1)-((2 x^2+4 x-3) (d/dx(2 x-1)))/(2 (2 x-1)^(3/2)) Differentiate the sum term by term and factor out constants: (2 (d/dx(x^2))+4 (d/dx(x))+d/dx(-3))/sqrt(2 x-1)-((2 x^2+4 x-3) (d/dx(-1)+2 (d/dx(x))))/(2 (2 x-1)^(3/2)) The derivative of -3 is zero: (2 (d/dx(x^2))+4 (d/dx(x))+0)/sqrt(2 x-1)-((2 x^2+4 x-3) (2 (d/dx(x))+d/dx(-1)))/(2 (2 x-1)^(3/2)) The derivative of -1 is zero: (2 (d/dx(x^2))+4 (d/dx(x)))/sqrt(2 x-1)-((2 x^2+4 x-3) (2 (d/dx(x))+0))/(2 (2 x-1)^(3/2)) The derivative of x is 1: (2 (d/dx(x^2))+1 4)/sqrt(2 x-1)-((2 x^2+4 x-3) (d/dx(x)))/(2 x-1)^(3/2) The derivative of x is 1: (2 (d/dx(x^2))+4)/sqrt(2 x-1)-(2 x^2+4 x-3)/(2 x-1)^(3/2) The derivative of x^2 is 2 x: Answer: | | (2 (2 x)+4)/sqrt(2 x-1)-(2 x^2+4 x-3)/(2 x-1)^(3/2)